# Graph Transformers Dream of Electric Flow

### v1

@@ Abstract
We show theoretically and empirically that the linear Transformer, when applied to graph data, can implement algorithms that solve canonical problems such as electric flow and eigenvector decomposition. The input to the Transformer is simply the graph incidence matrix; no other explicit positional encoding information is provided. We present explicit weight configurations for implementing each such graph algorithm, and we bound the errors of the constructed Transformers by the errors of the underlying algorithms. Our theoretical findings are corroborated by experiments on synthetic data. Additionally, on a real-world molecular regression task, we observe that the linear Transformer is capable of learning a more effective positional encoding than the default one based on Laplacian eigenvectors. Our work is an initial step towards elucidating the inner-workings of the Transformer for graph data.
@@ 1 Introduction
The Transformer architecture (Vaswani et al., 2017) has seen great success in the design of graph neural networks (GNNs) (Dwivedi and Bresson, 2020; Ying et al., 2021; Kim et al., 2022; Muller et al., 2023), and it has demonstrated impressive performance in many applications ranging from prediction of chemical properties to analyzing social networks (Yun et al., 2019; Dwivedi et al., 2023; Rong et al., 2020). In contrast to this empirical success, the underlying reasons for why Transformers adapt well to graph problems are less understood. Several works study the expressivity of attention-based architectures (Kreuzer et al., 2021; Kim et al., 2021, 2022; Ma et al., 2023), but such analyses often rely on the ability of neural networks to approximate arbitrary functions, and may require a prohibitive number of parameters.
This paper is motivated by the need to understand, at a mechanistic level, how the Transformer processes graph-structured data. Specifically, we study the ability of a linear Transformer to solve certain classes of graph problems. The linear Transformer is similar to the standard Transformer, but softmax-based activation is replaced with linear attention, and the MLP layers are replaced by linear layers. Notably, the linear Transformer contains no a priori knowledge of the graph structure; all information about the graph is provided via an incidence matrix $B$. For unweighted graphs, the columns of $B$ are just $\left\{-1,0,1\right\}$-valued indicator vectors that encode whether an edge touches a vertex; no other explicit positional or structural encodings are provided.
@@
Besides enhancing our understanding of Transformers, our results are also useful for the practical design of graph Transformers. In Sections 3 and 4, we show that essentially the same linear Transformer architecture is capable of learning a number of popular positional encodings (PE). In Section 6, we provide experimental evidence that the linear Transformer can learn better PEs than hard-coded PEs.
@@ 1.1 Summary of Contributions
Below, we summarize the main content of our paper.
@@
    In Section 5, we provide a more efficient, constrained version of the linear Transformer with much fewer parameters. Although the constrained Transformer is less expressive by definition, we show that it can nonetheless implement all the above-mentioned constructions. Further, we show in Lemma 10 that the parameter-efficient linear Transformer has desirable invariance and equivariance properties.
@@
    In Section 6, we demonstrate the advantage of using the linear Transformer as a replacement for Laplacian eigenvector positional encoding on a molecular regression benchmark using the QM9 and ZINC datasets (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014; Irwin et al., 2012). After replacing the Laplacian eigenvector-based PE with the linear Transformer, and training on the regression loss, we verify that the linear Transformer automatically learns a good PE for the downstream regression task that can outperform the original Laplacian eigenvector-based PE by a wide margin.
@@ 3 Heat Kernel ($e^{-s{\mathcal{L}}}$) Power Series $\log(1/\varepsilon)+s\lambda_{\max}$
Table 1: Summary of Error Bounds proved for various Transformer constructions. ${\mathcal{L}}$ is the graph Laplacian, ${\mathcal{L}}^{\dagger}$ denotes its pseudoinverse.
@@ 1.2 Related Work
Numerous authors have proposed different ways of adapting Transformers to graphs (Dwivedi and Bresson, 2020; Ying et al., 2021; Rampasek et al., 2022; Muller et al., 2023). A particularly promising approach is to use a positional encodings to incorporate structural information in the input, examples include Laplacian eigenvectors (Dwivedi and Bresson, 2020; Kreuzer et al., 2021), heat kernel (Choromanski et al., 2022), resistance distance and commute time (Ma et al., 2023; Velingker et al., 2024; Zhang et al., 2023) and shortest path distance (Ying et al., 2021). Lim et al. (2022) design a neural network to transform eigenvectors into an encoding that has certain invariance properties. Black et al. (2024) compare the expressivity of different PE schemes.
@@
Ma et al. (2023) propose a novel graph Transformer architecture that incorporates learned PEs based on learned relative random walk probabilities (RRWP). Kreuzer et al. (2021) also propose learning PEs parameterized by eigenfunctions. In comparison, using a linear-Transformer to learn PEs has a few advantages: First, self attention can implement operations such as orthogonalization that are necessary for learning eigenvectors (Lemma 6). Second, self-attention enables some highly-efficient algorithms based on multiplicative polynomial expansions (Lemma 4). For these reasons, linear Transformers can learn good PEs given only the incidence matrix as input.
Finally, we note several other relevant papers that are unrelated to graph neural networks. A recent body of work proves the ability of Transformers to implement learning algorithms, to explain the in-context learning phenomenon (Schlag et al., 2021; Von Oswald et al., 2023). Surprisingly, our construction for Lemma 1 bears several remarkable parallels to the gradient descent construction by Von Oswald et al. (2023). We conjecture that the proof of Lemma 4 may be applicable for understanding the GD++ algorithm in the same paper. In another direction, Charton (2021) show experimentally that Transformers can compute linear algebra operations such as eigenvector decomposition. Their work requires a relatively complicated matrix encoding and a large number of parameters.
@@ 2.1 Graphs
We use ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ to denote a graph with vertex set ${\mathcal{V}}$ and edge set ${\mathcal{E}}$; $n$ denotes $|{\mathcal{V}}|$ and $d$ represents $|{\mathcal{E}}|$. We often identify the vertex $v_{i}$ with its index $i$ for $i=1...n$, and the edge $e_{j}$ with $j$ for $j=1...d$. We will generally consider weighted graphs, where $r(\cdot):{\mathcal{E}}\to\mathbb{R}^{+}$ gives the edge weights. We will use $r_{j}:=r(e_{j})$. A central object of interest is the incidence matrix $B\in\mathbb{R}^{n\times d}$, defined as follows: to each edge $e_{j}\in{\mathcal{E}}$, assign an arbitrary orientation, i.e. $e=(u\to v)\in{\mathcal{E}}$. The matrix $B$ is given by
@@
Next, we define the graph Laplacian as ${\mathcal{L}}:=BB^{\top}\in\mathbb{R}^{n\times n}$. We will often need to refer to the maximum eigenvalue of ${\mathcal{L}}$, which we denote as $\lambda_{\max}$. Note that ${\mathcal{L}}$ always has 00 as its smallest eigenvalue, with corresponding eigenvector $\frac{1}{\sqrt{n}}\vec{1}$, the all-ones vector. This fact can be verified by noticing that each column of $B$ sums to 00. For a connected graph (as we will assume is the case throughout the paper), the second-smallest eigenvalue is always non-zero, and we will denote it as $\lambda_{\min}$.
Finally, we will frequently refer to the matrix $I_{n\times n}-\frac{1}{n}\vec{1}\vec{1}^{\top}$; for ease of reference, we will write $\hat{I}:=I_{n\times n}-\frac{1}{n}\vec{1}\vec{1}^{\top}$. This matrix is the projection onto $({\mathcal{L}})$, and it essentially functions as the identity matrix, as we work within $({\mathcal{L}})$ most of the time.
@@ 2.2 Linear Transformer
$\mathrm{Attn}_{W^{V},W^{Q},W^{K}}(Z):=W^{V}ZZ^{\top}{W^{Q}}^{\top}W^{K}Z.$ (2.2)
Unlike standard attention, (2.2) does not contain softmax activation. We construct an $L$-layer Transformer by stacking $L$ layers of the attention module (with linear feed-forward). To be precise, let $Z_{l}$ denote the output of the $(l-1)^{th}$ layer of the Transformer. Then
$Z_{l+1}:=Z_{l}+\mathrm{Attn}_{W^{V}_{l},W^{Q}_{l},W^{K}_{l}}{\left(Z_{l}\right)}+{W^{R}}Z_{l},$ (2.3)
where $W^{V}_{l},W^{Q}_{l},W^{K}_{l}$ are the value, query and key weight matrices of the linear attention module at layer $l$, and $W^{R}\in\mathbb{R}^{h\times h}$ is the weight matrix of the linear module. Henceforth, we let $W^{V}:=\{W^{V}_{l}\}_{l=0\ldots L}$, $W^{Q}:=\{W^{Q}_{l}\}_{l=0\ldots L}$, $W^{K}_{l}:=\{W^{K}_{l}\}_{l=0\ldots L}$, $W^{R}_{l}:=\{W^{R}_{l}\}_{l=0\ldots L}$ denote collections of the parameters across all layers of an $L$-layer Transformer.
Finally, we will often need to refer to specific rows of $Z_{l}$. We will use ${\left[Z_{l}\right]}_{i...j}$ to denote rows $i$ to $j$ of $Z_{l}$.
@@ 3 Transformers as powerful solvers for Laplacian problems
In this section, we discuss the capacity of the linear Transformer 2.3 to solve certain classes of canonical graph problems. We begin with the problem of Electric Flow in Section 3.2, where the constructed Transformer can be interpreted as implementing steps of gradient descent with respect to the energy of the induced flow. Subsequently, in Section 3.3, we provide constructions for computing the resistive embedding, as well as the heat kernel, based on implementing a truncated power series. Finally, in Section 3.4, we provide faster alternative constructions for solving electric flow and computing heat kernels, based on implementing a multiplicative polynomial expansion. In each case, we bound the error of the Transformer by the convergence rate of the underlying algorithms.
@@ 3.1 Additional Setup
We introduce some additional setup that is common to many lemmas in this section. We will generally consider an $L$-layer Transformer, as defined in (2.3), for some arbitrary $L\in\mathbb{Z}^{+}$. As in (2.3), we use $Z_{l}$ to denote the input to layer $l$. The input $Z_{0}$ will encode information about a graph ${\mathcal{G}}$, along with a number of demand vectors $\psi_{1}\ldots\psi_{k}\in\mathbb{R}^{n}$. We use $\Psi$ to denote the $n\times k$ matrix whose $i^{th}$ column is $\psi_{i}$. Unless otherwise stated, $Z_{l}\in\mathbb{R}^{(d+2k)\times n}$, where $n$ is the number of nodes, $d$ is the number of edges, and $k$ is the number of demands/queries. $Z_{0}^{\top}:=\begin{bmatrix}B,&\Psi,&0_{n\times k}\end{bmatrix}$.
On parameter size: In a straightforward implementation, the above Transformer has feature dimension $h=(d+2k)$. The size of $W^{Q},W^{K},W^{V},W^{R}$ is $O(h^{2})=O(d^{2}+k^{2})$, which is prohibitively large as $d$ can itself be $O(n^{2})$. The size of parameter matrices can be significantly reduced to $O(k^{2})$ by imposing certain constraints on the parameter matrices; we present this reduction in (5.1) in Section 5. For simplicity of exposition, lemmas in this section will use the naive implementation in (2.3). We verify later that all the constructions presented in this section can also be realized in (5.1).
@@ 3.2 Solving Electric Flow with Gradient Descent
(primal) $\min_{f\in\mathbb{R}^{d}}\quad\sum_{j=1}^{d}r_{j}f_{j}^{2}\qquad\qquad\text{subject to the constraint}\quad BR^{1/2}f=\psi.$ (3.1)
(dual) $-\min_{\phi\in\mathbb{R}^{n}}\phi^{\top}{\mathcal{L}}\phi-2\phi^{\top}\psi.$ (3.2)
The argument is standard; for completeness, we provide a proof of equivalence between (3.1) and (3.2) in (8.1) in Appendix 8.1. It follows that the optimizer $\phi^{*}$ of (3.2) has closed-form solution $\phi^{*}={\mathcal{L}}^{\dagger}\psi$. In Lemma 1 below, we show a simple construction that enables the Transformer in (2.3) to compute in parallel, the optimal potential assignments for a set of $k$ demands $\left\{\psi_{i}\right\}_{i=1...k}$, where $\psi_{i}\in\mathbb{R}^{n}$.
@@
An important practical motivation for solving the electric flow (or equivalently computing ${\mathcal{L}}^{\dagger}$) is to obtain the Effective Resistance matrix ${\mathcal{R}}\in\mathbb{R}^{n\times n}$. GNNs that use positional encodings derived from ${\mathcal{R}}$ have demonstrated state-of-the-art performance on numerous tasks, and can be shown to have good theoretical expressivity (Zhang et al., 2023; Velingker et al., 2024; Black et al., 2024). Formally, ${\mathcal{R}}$ is defined as ${\mathcal{R}}_{ij}:=(u_{i}-u_{j})^{\top}{\mathcal{L}}^{\dagger}(u_{i}-u_{j})$, where $u_{i}$ denotes the vector that has a $1$ in the $i^{th}$ coordinate, and 00s everywhere else. Intuitively, ${\mathcal{R}}_{ij}$ is the potential drop required to send 1-unit of electric flow from node $i$ to node $j$. Let $\ell\in\mathbb{R}^{n}$ denote the vector of diagonals of ${\mathcal{L}}^{\dagger}$ (i.e. $\ell_{i}:={\mathcal{L}}_{ii}$). Then ${\mathcal{R}}=\vec{1}\ell^{\top}+\ell\vec{1}^{\top}-2{\mathcal{L}}^{\dagger}$; thus computing ${\mathcal{L}}^{\dagger}$ essentially also computes ${\mathcal{R}}$.
@@
Lemma 1 (Transformer solves Electric Flow by implementing Gradient Descent).
@@
Consider the setup in Section 3.1. Assume that $\langle{\psi_{i},\vec{1}}\rangle=0$ for each $i=1...k$. For any $\delta>0$ and for any $L$-layer Transformer, there exists a choice of weights $W^{V},W^{Q},W^{K},W^{R}$, such that each layer of the Transformer (2.3) implements a step of gradient descent with respect to the dual electric flow objective in (3.2), with stepsize $\delta$. Consequently, the following holds for all $i=1...k$ and for any graph Laplacian with maximum eigenvalue $\lambda_{\max}\leq 1/\delta$ and minimum nontrivial eigenvalue $\lambda_{\min}$:
@@
${\left\|{\left[Z_{L}\right]}_{d+k+i}^{\top}-{\mathcal{L}}^{\dagger}\psi_{i}\right\|}_{2}\leq\frac{\exp{\left(-\delta\lambda_{min}/2\right)}}{\sqrt{\lambda_{min}}}\|\psi_{i}\|_{2}^{2}{\left\|\psi_{i}\right\|}_{2}.$
@@
Discussion. In the special case when $k=n$ and $\Psi:=I_{n\times n}-\frac{1}{n}\vec{1}\vec{1}^{\top}$, ${\left[Z_{L}\right]}_{d+k+1...d+2k}\approx{\mathcal{L}}^{\dagger}$. An $\varepsilon$-approximate solution requires $L=O(\log(1/\varepsilon))$ layers; this is exactly the convergence rate of gradient descent on a Lipschitz smooth and strongly convex function. In Lemma 4 below, we show an alternate construction that reduces this to $O(\log\log(1/\varepsilon))$.
@@
We defer details of the proof of Lemma 1 to Appendix 8.1, and we provide the explicit choice of weight matrices in (8.3). The constructed weight matrices are very sparse; each of $W^{V},W^{Q},W^{K},W^{R}$ contains a single identity matrix in a sub-block. This sparse structure makes it possible to drastically reduce the number of parameters needed, which we exploit in Section 5 to design a more parameter efficient Transformer. We provide experimental validation of Lemma 1 in Section 3.5.

### v2

@@ Abstract
We show theoretically and empirically that the linear Transformer, when applied to graph data, can implement algorithms that solve canonical problems such as electric flow and eigenvector decomposition. The Transformer has access to information on the input graph only via the graph’s incidence matrix. We present explicit weight configurations for implementing each algorithm, and we bound the constructed Transformers’ errors by the errors of the underlying algorithms. Our theoretical findings are corroborated by experiments on synthetic data. Additionally, on a real-world molecular regression task, we observe that the linear Transformer is capable of learning a more effective positional encoding than the default one based on Laplacian eigenvectors. Our work is an initial step towards elucidating the inner-workings of the Transformer for graph data. Code is available at https://github.com/chengxiang/LinearGraphTransformer
@@ 1 Introduction
The Transformer architecture (Vaswani et al., 2017) has seen great success in the design of graph neural networks (GNNs) (Dwivedi & Bresson, 2020; Ying et al., 2021; Kim et al., 2022; Müller et al., 2023), and it has demonstrated impressive performance in many applications ranging from prediction of chemical properties to analyzing social networks (Yun et al., 2019; Dwivedi et al., 2023; Rong et al., 2020). In contrast to this empirical success, the underlying reasons for why Transformers adapt well to graph problems are less understood. Several works study the expressivity of attention-based architectures (Kreuzer et al., 2021; Kim et al., 2021; 2022; Ma et al., 2023), but such analyses often rely on the ability of neural networks to approximate arbitrary functions, and may require a prohibitive number of parameters.
This paper is motivated by the need to understand, at a mechanistic level, how the Transformer processes graph-structured data. Specifically, we study the ability of a linear Transformer to solve certain classes of graph problems. The linear Transformer is similar to the standard Transformer, but softmax-based activation is replaced with linear attention, and MLP layers are replaced by linear layers. Notably, the linear Transformer contains no a priori knowledge of the graph structure; all information about the graph is provided via an incidence matrix $B$. For unweighted graphs, the columns of $B$ are just $\left\{-1,0,1\right\}$-valued indicator vectors that encode whether an edge touches a vertex; no other explicit positional or structural encodings are provided.
@@
Besides enhancing our understanding of Transformers, our results are also useful for the practical design of graph Transformers. In Sections 3 and 4, we show that the same linear Transformer architecture is capable of learning a number of popular positional encodings (PE). In Section 6, we provide experimental evidence that the linear Transformer can learn better PEs than hard-coded PEs.
@@ 1.1 Summary of Contributions
Below, we summarize the main contributions of our paper.
@@
    In Section 5, we propose a more efficient version of the linear Transformer with much fewer parameters. We show that this more efficient linear Transformer can nonetheless implement all the above-mentioned constructions. Further, we show in Lemma 10 that the parameter-efficient linear Transformer has desirable invariance and equivariance properties.
@@
    In Section 6, we demonstrate the advantage of using the linear Transformer as a replacement for Laplacian eigenvector PE on a molecular regression benchmark using the QM9 and ZINC datasets (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014; Irwin et al., 2012). After replacing the Laplacian eigenvector-based PE with the linear Transformer, and training on the regression loss, we verify that the linear Transformer automatically learns a good PE for the downstream regression task that can outperform the original Laplacian eigenvector PE by a wide margin.
@@ 3 Heat Kernel ($e^{-s{\mathcal{L}}}$) Power Series $\log(1/\varepsilon)+s\lambda_{\max}$
Table 1: Summary of Error Bounds proved for various Transformer constructions. ${\mathcal{L}}$ is the graph Laplacian; ${\mathcal{L}}^{\dagger}$ denotes its pseudoinverse.
@@ 1.2 Related Work
Numerous authors have proposed different ways of adapting Transformers to graphs (Dwivedi & Bresson, 2020; Ying et al., 2021; Rampávsek et al., 2022; Müller et al., 2023). A particularly promising approach is to use a suitable positional encoding to incorporate structural information in the input, examples include Laplacian eigenvectors (Dwivedi & Bresson, 2020; Kreuzer et al., 2021), heat kernel (Choromanski et al., 2022), resistance distance and commute time (Ma et al., 2023; Velingker et al., 2024; Zhang et al., 2023) and shortest path distance (Ying et al., 2021). Lim et al. (2022) designed a neural network to transform eigenvectors into an encoding that has certain invariance properties. Black et al. (2024) compared the expressivity of different PE schemes. (Srinivasan & Ribeiro, 2020) studied the relationship between PE and structural graph representations.
@@
A number of works have explored the use of learned PEs (Mialon et al., 2021; Eliasof et al., 2023; Park et al., 2022; Ma et al., 2023; Kreuzer et al., 2021; Dwivedi et al., 2021). In particular Ma et al. (2023) is based on relative random walk probabilities (RRWP), and Kreuzer et al. (2021)’s approach is based on learned eigenfunctions. In comparison, using a linear-Transformer to learn PEs has two advantages: 1. self attention can implement operations such as orthogonalization, necessary for learning multiple eigenvectors (Lemma 6). Second, self-attention enables highly-efficient algorithms based on multiplicative polynomial expansions (Lemma 4). These enable linear Transformers to learn good PEs given only the incidence matrix as input.
Finally, we note several relevant papers that are unrelated to graph neural networks. A recent body of work proves the ability of Transformers to implement learning algorithms, to explain the in-context learning phenomenon (Schlag et al., 2021; Von Oswald et al., 2023). Surprisingly, our construction for Lemma 1 bears several remarkable parallels to the gradient descent construction by Von Oswald et al. (2023). We conjecture that the proof of Lemma 4 may be applicable for understanding the GD++ algorithm in the same paper. In another direction, Charton (2021) show experimentally that Transformers can compute linear algebra operations such as eigenvector decomposition. Their work requires a relatively complicated matrix encoding and a large number of parameters.
@@ 2.1 Graphs
We use ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ to denote a graph with vertex set ${\mathcal{V}}$ and edge set ${\mathcal{E}}$; let $n:=|{\mathcal{V}}|$ and $d:=|{\mathcal{E}}|$. We often identify the vertex $v_{i}$ with its index $i$ for $i=1...n$, and the edge $e_{j}$ with $j$ for $j=1...d$. In general, ${\mathcal{G}}$ has weighted edges, defined by $r(\cdot):{\mathcal{E}}\to\mathbb{R}^{+}$. We let $r_{j}:=r(e_{j})$. A central object of interest is the incidence matrix $B\in\mathbb{R}^{n\times d}$, defined as follows: to each edge $e_{j}=(u\sim v)\in{\mathcal{E}}$, assign an arbitrary orientation, e.g. $e_{j}=(u\to v)$. The incidence matrix $B$ is given by
@@
We define the graph Laplacian as ${\mathcal{L}}:=BB^{\top}\in\mathbb{R}^{n\times n}$. We use $\lambda_{\max}$ to denote the maximum eigenvalue of ${\mathcal{L}}$. Note that ${\mathcal{L}}$ always has 00 as its smallest eigenvalue, with corresponding eigenvector $\frac{1}{\sqrt{n}}\vec{1}$, the all-ones vector. This fact can be verified by noticing that each column of $B$ sums to 00. For a connected graph (as we will assume is the case throughout the paper), the second-smallest eigenvalue is always non-zero, and we will denote it as $\lambda_{\min}$. Finally, we define $\hat{I}_{n\times n}:=I_{n\times n}-\frac{1}{n}\vec{1}\vec{1}^{\top}$. $\hat{I}_{n\times n}$ is the projection onto $({\mathcal{L}})$, and functions as the identity matrix when working within $({\mathcal{L}})$.
@@ 2.2 Linear Transformer
$\mathrm{Attn}_{W^{V},W^{Q},W^{K}}(Z):=W^{V}ZZ^{\top}{W^{Q}}^{\top}W^{K}Z.$ (2)
Unlike standard attention, (2) does not contain softmax activation. We construct an $L$-layer Transformer by stacking $L$ layers of the attention module (augmented with a linear operation on the skip connection). To be precise, let $Z_{l}$ denote the output of the $(l-1)^{th}$ layer of the Transformer. Then
$Z_{l+1}:=Z_{l}+\mathrm{Attn}_{W^{V}_{l},W^{Q}_{l},W^{K}_{l}}{\left(Z_{l}\right)}+{W^{R}}Z_{l},$ (3)
where $W^{V}_{l},W^{Q}_{l},W^{K}_{l}$ are the value, query and key weight matrices of the linear attention module at layer $l$, and $W^{R}\in\mathbb{R}^{h\times h}$ is the weight matrix of the linear operation. Henceforth, we let $W^{V}:=\{W^{V}_{l}\}_{l=0\ldots L}$, $W^{Q}:=\{W^{Q}_{l}\}_{l=0\ldots L}$, $W^{K}_{l}:=\{W^{K}_{l}\}_{l=0\ldots L}$, $W^{R}_{l}:=\{W^{R}_{l}\}_{l=0\ldots L}$ denote collections of the parameters across all layers of an $L$-layer Transformer. Finally, we will often need to refer to specific rows of $Z_{l}$. We will use ${\left[Z_{l}\right]}_{i...j}$ to denote rows $i$ to $j$ of $Z_{l}$.
@@ 3 Transformers as powerful solvers for Laplacian problems
In this section, we discuss the capacity of the linear Transformer 3 to solve certain classes of canonical graph problems. We begin with the problem of Electric Flow in Section 3.2, where the constructed Transformer can be interpreted as implementing steps of gradient descent with respect to the energy of the induced flow. Subsequently, in Section 3.3, we provide constructions for computing the resistive embedding, as well as the heat kernel, based on implementing a truncated power series. Finally, in Section 3.4, we provide faster alternative constructions for solving electric flow and computing heat kernels, based on implementing a multiplicative polynomial expansion. In each case, we bound the error of the Transformer by the convergence rate of the underlying algorithms.
@@ 3.1 Additional Setup
We introduce some additional setup that is common to many lemmas in this section. We will generally consider an $L$-layer Transformer, as defined in (3), for some arbitrary $L\in\mathbb{Z}^{+}$. As in (3), we use $Z_{l}$ to denote the input to layer $l$. The input $Z_{0}$ will encode information about a graph ${\mathcal{G}}$, along with a number of demand vectors $\psi_{1}\ldots\psi_{k}\in\mathbb{R}^{n}$. We use $\Psi$ to denote the $n\times k$ matrix whose $i^{th}$ column is $\psi_{i}$. Unless otherwise stated, $Z_{l}\in\mathbb{R}^{(d+2k)\times n}$, where $n$ is the number of nodes, $d$ is the number of edges, and $k$ is the number of demands/queries. $Z_{0}^{\top}:=\begin{bmatrix}B,&\Psi,&0_{n\times k}\end{bmatrix}$.
On parameter size: In a straightforward implementation, the above Transformer has feature dimension $h=(d+2k)$. The size of $W^{Q},W^{K},W^{V},W^{R}$ is $O(h^{2})=O(d^{2}+k^{2})$, which is prohibitively large as $d$ can itself be $O(n^{2})$. The size of parameter matrices can be significantly reduced to $O(k^{2})$ by imposing certain constraints on the parameter matrices; we present this reduction in (7) in Section 5. For simplicity of exposition, lemmas in this section will use the naive implementation in (3). We verify later that all the constructions presented in this section can also be realized in (7).
@@ 3.2 Solving Electric Flow with Gradient Descent
(primal) $\min_{f\in\mathbb{R}^{d}}\quad\sum_{j=1}^{d}r_{j}f_{j}^{2}\qquad\qquad\text{subject to the constraint}\quad BR^{1/2}f=\psi.$ (4)
(dual) $-\min_{\phi\in\mathbb{R}^{n}}\phi^{\top}{\mathcal{L}}\phi-2\phi^{\top}\psi.$ (5)
The argument is standard; for completeness, we provide a proof of equivalence between (4) and (5) in (8) in Appendix 10.1. It follows that the optimizer $\phi^{*}$ of (5) has closed-form solution $\phi^{*}={\mathcal{L}}^{\dagger}\psi$. In Lemma 1 below, we show a simple construction that enables the Transformer in (3) to compute in parallel, the optimal potential assignments for a set of $k$ demands $\left\{\psi_{i}\right\}_{i=1...k}$, where $\psi_{i}\in\mathbb{R}^{n}$.
@@
An important practical motivation for solving the electric flow (or equivalently computing ${\mathcal{L}}^{\dagger}$) is to obtain the Effective Resistance matrix ${\mathcal{R}}\in\mathbb{R}^{n\times n}$. GNNs that use positional encodings derived from ${\mathcal{R}}$ have demonstrated state-of-the-art performance on numerous tasks, and can be shown to have good theoretical expressivity (Zhang et al., 2023; Velingker et al., 2024; Black et al., 2024).
Formally, ${\mathcal{R}}$ is defined as ${\mathcal{R}}_{ij}:=(u_{i}-u_{j})^{\top}{\mathcal{L}}^{\dagger}(u_{i}-u_{j})$, where $u_{i}$ denotes the vector that has a $1$ in the $i^{th}$ coordinate, and 00s everywhere else. Intuitively, ${\mathcal{R}}_{ij}$ is the potential drop required to send 1-unit of electric flow from node $i$ to node $j$. Let $\ell\in\mathbb{R}^{n}$ denote the vector of diagonals of ${\mathcal{L}}^{\dagger}$ (i.e. $\ell_{i}:={\mathcal{L}}_{ii}$). Then ${\mathcal{R}}=\vec{1}\ell^{\top}+\ell\vec{1}^{\top}-2{\mathcal{L}}^{\dagger}$; thus computing ${\mathcal{L}}^{\dagger}$ essentially also computes ${\mathcal{R}}$.
@@
Lemma 1 (Transformer solves Electric Flow by implementing Gradient Descent)
@@
Consider the setup in Section 3.1. Assume that $\langle{\psi_{i},\vec{1}}\rangle=0$ for each $i=1...k$. For any $\delta>0$ and for any $L$-layer Transformer, there exists a choice of weights $W^{V},W^{Q},W^{K},W^{R}$, such that each layer of the Transformer (3) implements a step of gradient descent with respect to the dual electric flow objective in (5), with stepsize $\delta$. Consequently, the following holds for all $i=1...k$ and for any graph Laplacian with maximum eigenvalue $\lambda_{\max}\leq 1/\delta$ and minimum nontrivial eigenvalue $\lambda_{\min}$:
@@
${\left\|{\left[Z_{L}\right]}_{d+k+i}^{\top}-{\mathcal{L}}^{\dagger}\psi_{i}\right\|}_{2}\leq\frac{\exp{\left(-\delta L\lambda_{\min}/2\right)}}{\sqrt{\lambda_{\min}}}{\left\|\psi_{i}\right\|}_{2}.$
@@
Discussion. In the special case when $k=n$ and $\Psi:=I_{n\times n}-\frac{1}{n}\vec{1}\vec{1}^{\top}$, ${\left[Z_{L}\right]}_{d+k+1...d+2k}\approx{\mathcal{L}}^{\dagger}$. An $\varepsilon$-approximate solution requires $L=O(\log(1/\varepsilon))$ layers; this is exactly the convergence rate of gradient descent on a Lipschitz smooth and strongly convex function. We defer details of the proof of Lemma 1 to Appendix 10.1, and we provide the explicit choice of weight matrices in (10). On a high level, we show that a single attention layer can implement exactly a single gradient descent step wrt the dual objective $F_{\psi}$ (5), i.e. ${\left[Z_{{\color[rgb]{0,0,1}l+1}}\right]}_{d+k+i}^{\top}:={\left[Z_{{\color[rgb]{0,0,1}l}}\right]}_{d+k+i}^{\top}-\delta\nabla F_{\psi_{i}}({\left[Z_{{\color[rgb]{0,0,1}l}}\right]}_{d+k+i}^{\top}).$ The $\log(1/\varepsilon)$ rate then follows immediately from the convergence rate of gradient descent, combined with smoothness and (restricted) strong convex of $F_{\psi}$. In Lemma 4 below, we show an alternate construction that reduces this to $O(\log\log(1/\varepsilon))$.
@@
We highlight that the constructed weight matrices are very sparse; each of $W^{V},W^{Q},W^{K},W^{R}$ contains a single identity matrix in a sub-block. This sparse structure makes it possible to drastically reduce the number of parameters needed, which we exploit in Section 5 to design a more parameter efficient Transformer. We provide experimental validation of Lemma 1 in Section 3.5.